Find the general equation of the line: y = (2)/(3) x - 7
a)2x-3y=21
b)2x+3y=21
c)2x-3y=-21
d)2x+3y=-21

Answer by rwm(914)

multiply by 3
3y=2x-21
subtract 2x
3y-2x=-21
not D) nor C)
3y=2x-21
add 21
3y+21=2x
subtract 3y
2x-3y=21
A)